OpenStudy (anonymous):

What is the geometric mean of 5 and 15?

mathslover (mathslover):

geometric mean = \(\large{(a_1*a_2*..*a_n)^\frac{1}{n}}\) where a1 is your first number, a2 is your second number, and so forth n is the number of entries

mathslover (mathslover):

so we have : a1 = 5, a2 = 15 and n = 2 so :geometric mean =\(\large{\sqrt{ (5*15)}}\)

mathslover (mathslover):

can u do that now?

mathslover (mathslover):

\[\large{\sqrt{5*15}=\sqrt{5*5*3}}\]

OpenStudy (anonymous):

thats the same way i did it but my choices are only 1.\[5\sqrt{3}\] 2.75 3.\[3\sqrt{5}\] 4.37.5

OpenStudy (anonymous):

hello?

mathslover (mathslover):

1. is correct

mathslover (mathslover):

sorry for late reply

mathslover (mathslover):

\[\large{5\sqrt{3}}\]